A rectangle has one side that is 13 inches long and another side that is 1 foot long. Find the perimeter of the rectangle. Write your answer as a fraction using feet. I tried converting 13 inches to 1 foot, 1 inch, but I don't know how to put it as a fraction and if I actually did the right method.

Answer by mananth(16946)

one side is 1 foot
other is 1foot +1 foot + 1 inch
= 2 1/12 = 25/12 feet
Perimeter = 2*( 1+(25/12))
=2*((12+25)/12)
=2*(37/12)
= 37/6
=6 1/6 feet